Justin is 4 years older than Ethan. In 9 years the sum of their ages will be 34. Find the age of each boy now.
Answer by stanbon(75887)

Justin is 4 years older than Ethan. In 9 years the sum of their ages will be 34. Find the age of each boy now.
------
Equations:
j = e + 4
(j+9)+(e+9) = 34
---
Substitute for "j" and solve for "e":
(e+4+9) + (e+9) = 34
---
2e + 22 = 34
2e = 12
e = 6 (Ethan's age now)
-----
Solve for "j":
j = e+4
j = 6+4
j = 10 (Justin's age now)
=========
Cheers,
Stan H.
========